# Efficient Action Counting with Dynamic Queries

### 03 Mar 2024 v1

@@ 1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications ranging from pedestrian detection [29] to fitness monitoring [11].
Pioneer methods  [16, 1, 7, 38, 27, 37, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [17, 9, 13, 46] and show promising performances. Notably, most of these methods, such as RepNet [9] and TransRAC [13], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig. 1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
@@ 2.1 Temporal Repetition Counting
Traditional methods [16, 1, 7, 38, 27, 37, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [17, 9, 13, 46] have demonstrated remarkable performance improvements. Notably, RepNet [9] and TransRAC [13] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [46] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
@@ 4.1 Datasets and Metrics
UCFRep dataset [46] is a subset of the UCF101 dataset [35], including fitness videos and daily life videos. Following previous work [13], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
@@ 4.3 Comparison to State-of-the-arts
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [13] dataset. We compare with the SOTA action recognition [10, 24, 23], action segmentation [14], and TRC [9, 46, 13] approaches as explored in TransRAC [13]. We further adapt recent DETR-style action detection approaches [22, 44, 33] for the TRC task. We change their output layers accordingly and train them on RepCountA.
Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset. We further calculate MAE and OBO metrics for short-, medium-, and long-period test actions. normal-$\dagger$ denotes action detection approaches.
@@
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 13] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [22, 44, 33] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not provide them with the ability to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [13], we evaluate the model trained on the RepCountA dataset [13] on UCFRep [46] test set. Tab. 2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
Table 2: Generalization comparison on UCFRep [46] dataset. We calculate MAE and OBO metrics for short-, medium-, and long-period test actions.
@@
Table 3: Comparison of computational complexity.
@@ 4.4 Qualitative Results
Figure 3: RepCountA visualization results. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
@@
Figure 5: UCFRep visualization results. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [46].

### 05 Mar 2024 v2

@@ 1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [48].
Pioneer methods  [16, 1, 7, 38, 28, 37, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [17, 9, 13, 46, 18] and show promising performances. Notably, most of these methods, such as RepNet [9] and TransRAC [13], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig. 1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
@@ 2.1 Temporal Repetition Counting
Traditional methods [16, 1, 7, 38, 28, 37, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [17, 9, 13, 46, 18] have demonstrated remarkable performance improvements. Notably, RepNet [9] and TransRAC [13] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [46] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
@@ 4.1 Datasets and Metrics
UCFRep dataset [46] is a subset of the UCF101 dataset [35], including fitness videos and daily life videos. Following previous work [13, 18], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
@@ 4.3 Comparison to State-of-the-arts
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [13] dataset following previous work [13, 18] in Tab. 1. We compare with the SOTA action recognition [10, 25, 24], action segmentation [14] methods (top block), and TRC [9, 46, 13, 18] approaches (second block). We further adapt recent DETR-style action detection approaches [23, 44, 33] for the TRC task (third block). We change their output layers accordingly and train them on RepCountA.
Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with normal-$\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
@@
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 13, 18] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [23, 44, 33] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not provide them with the ability to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [13, 18], we evaluate the model trained on the RepCountA dataset [13] on UCFRep [46] test set. Tab. 2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
Table 2: Generalization comparison with SOTA TRC methods on UCFRep [46] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
@@
Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
@@ 4.4 Qualitative Results
Figure 3: Qualitative results on RepCountA dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
@@
Figure 5: Qualitative results on UCFRep dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [46].